# Seat Arrangement Problems under B-utility and W-utility

### v1

@@ 4 Envy-freeness (EFA)
We first show that in an envy-free arrangement $\sigma$ agent $p_{i}$ is allocated to the centre vertex in the $i$-th star (or a star with the same number of vertices). We have that every star has at least four leaves. However, every agent that is not of type $p_{x}$ has a preference of $1$ for at most three other agents. Therefore, if such an agent $a$ is assigned to a centre vertex, there exists at least one agent $b$ assigned to a leaf vertex such that $f_{a}(b)=f_{b}(a)=0$, a contradiction. Furthermore, $p_{i}$ cannot be assigned to the centre vertex in a star with more vertices than the $i$-th star. Assume otherwise and let $p_{i}$ be assigned to the centre vertex of the $j$-th, such that the number of vertices of this star is larger than the number of vertices of the $i$-th star. Thus, $|Q_{j}|+|R_{j}|-1>|Q_{i}|+|R_{i}|-1$. The way the sets are constructed, this inequality implies that $|Q_{j}|>|Q_{i}|$ and $|R_{j}|>|R_{i}|$, meaning the $j$-th star has at least two more vertices than the $i$-th star, and that $|Q_{j}|+|R_{j}|-1>|Q_{i}|+|R_{i}|$. Hence, if $p_{i}$ is assigned to the centre of the $j$-th star, then there exists at least one agent $a$ assigned to a leaf vertex that does not belong to $Q_{i}\cup R_{i}$, so $f_{p_{i}}(a)=f_{a}(p_{i})=0$, a contradiction.
As $p_{i}$ must be assigned to the centre vertex of the $i$-th star, and all agents in $R_{i}$ have a preference of $1$ over $p_{i}$ and of 00 over everyone else, these agents must be assigned to leaves in the $i$-th star. The remaining leaves must be occupied by agents from $Q_{i}$, as any other agent would have utility 00. Hence, exactly one agent from $Q_{i}$ is not assigned to a vertex from the $i$-th star, or any other star. Let this agent be $q_{i}^{j,k}$. In order for its utility to be $1$ and thus not be envious of a neighbour of $p_{i}$, one of the two other vertices in the cycle it has been assigned to must be occupied by either $q_{j}^{i,k}$ or $q_{k}^{i,j}$. Without loss of generality, let this be $q_{j}^{i,k}$. If the third vertex is occupied by some agent $a\neq q_{k}^{i,j}$, its utility would be 00. Hence, the third vertex must be assigned to $q_{k}^{i,j}$. That is, we have a triangle in the original instance with vertices $v_{i},v_{j},v_{k}$. As only one agent from $Q_{i}$ is in some cycle component, we form only one triangle with vertex $v_{i}$, meaning that we obtain a partition in $I$. ∎
@@
If preferences are symmetric and strict, instead of binary and symmetric, EFA with B-utility is still NP-hard. We will prove that later, in Theorem 8. Likewise, finding an envy-free arrangement is NP-hard when the Seat Arrangement instance $I$ has utility type S-utility. By Bodlaender et al. [4], this result holds even for symmetric preferences, and by Massand and Simon [16] the result holds even for binary preferences. We strengthen them, and show that it holds even for symmetric and binary, or symmetric and strict preferences.
@@
Let $I=(P,F,G,S)$ be an instance of Seat Arrangement with a planar seat graph of bounded degree $G$, symmetric and binary agents’ preferences, and S-utility. Then it is NP-complete to decide whether there exists an envy-free arrangement. NP-completeness also holds if agents’ preferences are symmetric and strict.
@@
For both restrictions, we will adapt the reduction from Theorem 3. For the symmetric and binary case, the agent set $P$, the seat graph $H$, and preference functions are identical. If $G$ has a partition into triangles, we build the same envy-free assignment from the proof of Theorem 3. Conversely, if the Seat Arrangement instance $I$ has an envy-free arrangement $\sigma$, every agent has utility at least $1$, mirroring Claim 1 from the proof of Theorem 3, and Claim 2 from the proof of Theorem 3 still applies, with minor modifications to its proof.
In the case of the symmetric and strict preferences, the agent set and seat graph are still the same. Regarding preferences, we start with the preference functions from the proof of Theorem 3. Then we break ties arbitrarily, using infinitesimally small positive constants. For example, if for agents $a,b,c$ $f_{a}(b)=f_{b}(a)=0$, $f_{a}(c)=f_{c}(a)=0$, and $f_{b}(c)=f_{c}(b)=0$, now $f_{a}(b)=f_{b}(a)=\varepsilon_{1}$, $f_{a}(c)=f_{c}(a)=\varepsilon_{2}$, and $f_{b}(c)=f_{c}(b)=\varepsilon_{3}$, where $\varepsilon_{1},\varepsilon_{2},\varepsilon_{3}$ are arbitrarily close to 00, positive, and pairwise distinct. Just like in the previous case, if $G$ can be partitioned into triangles we can build an envy-free arrangement, and if there is an envy-free arrangement every agent has utility at least $1$, as per Claim 1 from the proof of Theorem 3, and Claim 2 from the same proof still applies, with minor modifications to its proof. ∎
Similarly, in an instance with W-utility, finding an envy-free arrangement or reporting that none exists is NP-hard, even with symmetric and binary preferences, or symmetric and strict preferences, as we show with the following results. Their proofs are included in Appendix A.
@@
Let $I=(P,F,G,W)$ be an instance of Seat Arrangement with a planar seat graph $G$ with maximum degree $2$, symmetric and binary preferences, and W-utility. Then it is NP-complete to decide whether there exists an envy-free arrangement.
@@
The proof is included in Appendix A. ∎
@@
Let $I=(P,F,G,W)$ be an instance of Seat Arrangement with a planar seat graph of bounded degree $G$, symmetric and strict preferences, and W-utility. Then it is NP-complete to decide whether there exists an envy-free arrangement.

### v2

@@ 4 Envy-freeness (EFA)
We first show that in an envy-free arrangement $\sigma$ agent $p_{i}$ is allocated to the centre vertex in the $i$-th star (or a star with the same number of vertices). We have that every star has at least four leaves. However, every agent that is not of type $p_{x}$ has a preference of $1$ for at most three other agents. Therefore, if such an agent $a$ is assigned to a centre vertex, there exists at least one agent $b$ assigned to a leaf vertex such that $f_{a}(b)=f_{b}(a)=0$, a contradiction. Furthermore, $p_{i}$ cannot be assigned to the centre vertex in a star with more vertices than the $i$-th star. Assume otherwise and let $p_{i}$ be assigned to the centre vertex of the $j$-th star, such that the number of vertices of this star is larger than the number of vertices of the $i$-th star. Thus, $|Q_{j}|+|R_{j}|-1>|Q_{i}|+|R_{i}|-1$. The way the sets are constructed, this inequality implies that $|Q_{j}|>|Q_{i}|$ and $|R_{j}|>|R_{i}|$, meaning the $j$-th star has at least two more vertices than the $i$-th star, and that $|Q_{j}|+|R_{j}|-1>|Q_{i}|+|R_{i}|$. Hence, if $p_{i}$ is assigned to the centre of the $j$-th star, then there exists at least one agent $a$ assigned to a leaf vertex that does not belong to $Q_{i}\cup R_{i}$, so $f_{p_{i}}(a)=f_{a}(p_{i})=0$, a contradiction. Similarly, $p_{i}$ cannot be assigned to the centre vertex in a star with fewer vertices than the $i$-th star, because this implies that some other $p_{j}$ is assigned to the centre vertex of a star with more vertices than the $j$-th, which we have seen would imply that $u_{\sigma}(a)=0$ for some $a\in P$, a contradiction.
As $p_{i}$ must be assigned to the centre vertex of the $i$-th star, and all agents in $R_{i}$ have a preference of $1$ over $p_{i}$ and of 0 over everyone else, these agents must be assigned to leaves in the $i$-th star. The remaining leaves must be occupied by agents from $Q_{i}$, as any other agent would have utility 0. Hence, exactly one agent from $Q_{i}$ is not assigned to a vertex from the $i$-th star, or any other star. Let this agent be $q_{i}^{j,k}$. In order for its utility to be $1$ and thus not be envious of a neighbour of $p_{i}$, one of the two other vertices in the cycle it has been assigned to must be occupied by either $q_{j}^{i,k}$ or $q_{k}^{i,j}$. Without loss of generality, let this be $q_{j}^{i,k}$. If the third vertex is occupied by some agent $a\neq q_{k}^{i,j}$, its utility would be 0. Hence, the third vertex must be assigned to $q_{k}^{i,j}$. That is, we have a triangle in the original instance with vertices $v_{i},v_{j},v_{k}$. As only one agent from $Q_{i}$ is in some cycle component, we form only one triangle with vertex $v_{i}$, meaning that we obtain a partition in $I$. ∎
@@
If preferences are symmetric and strict, instead of binary and symmetric, EFA with B-utility is still NP-hard. In fact, even with $1$-dimensional and strict preferences, it is NP-complete to determine the existence of an envy-free arrangement. In order to prove this statement, we will reduce from the Bin Packing problem [garey_computers_1979]. In an instance of Bin Packing, we are given a set $X$ of items, such that each item $x_{i}\in X$ has a size $s_{i}\in\mathbb{N}$, a capacity $C\in\mathbb{N}$ for each bin, and a number of bins $K$. The set of all item sizes is $S$. The goal is to decide whether $X$ can be partitioned into at most $K$ different subsets such that the sum of the sizes of the elements in each subset is not greater than $C$.
Given an instance of Bin Packing, if $\sum_{x_{i}\in X}s_{i}<C\cdot K$, then we will add $C\cdot K-\sum_{x_{i}\in X}s_{i}$ elements of size $1$ each. Clearly this modified instance has a bin packing if and only if the previous instance has a bin packing too, because we can use the extra elements to “fill” any bin that is not full. Then, in order to ensure that the size of all elements is at least $2$, we will multiply every size by $2$, as well as the bin capacity $C$. These two changes are done to ensure that the transformed seat arrangement instance is a valid instance where the number of agents equals the number of vertices, and to ensure that from each item $x_{i}\in X$ we can obtain at least $2$ agents.
@@
Let $I=(P,F,G,B)$ be a seat arrangement instance with B-utility, $1$-dimensional strict preferences with unique positions, and a seat graph in which every connected component is a path with the same number of vertices. Then Envy-Free Arrangement is NP-hard.
@@
Let $(X,S,C,K)$ be an instance of Bin Packing, where $X$ is the set of elements, $C$ the bin capacity, and $K$ the target number of bins. The minimum size of every item is $2$, and $\sum_{x_{i}\in X}s_{i}=C\cdot K$. The bins in the Bin Packing will be labelled $C_{1},\dots,C_{K}$. Construct the seat arrangement instance $I=(P,F,G,B)$ as follows. Take item $x_{1}\in X$ and create $s_{1}$ agents, $p_{1}^{j}\in P$ ($1\leq j\leq s_{1}$). All of these agents belong to subset $P_{1}\subset P$. Lay them out in the real line starting with $p_{1}^{1}$, which will be placed at 0. Then $p_{1}^{2}$ will be placed a distance of exactly $1$ to the right of $p_{1}^{1}$, $p_{1}^{3}$ will be placed a distance of $1+\varepsilon$ to the right of $p_{1}^{2}$, where $\varepsilon$ is a suitable small positive number, $p_{1}^{4}$ will be placed a distance of $1+2\cdot\varepsilon$ to the right of $p_{1}^{3}$, and so on, until $p_{1}^{s_{1}}$, which will be placed a distance of $1+(s_{1}-2)\cdot\varepsilon$ to the right of $p_{1}^{s_{1}-1}$. After completing the placement of all agents associated with item $x_{1}$, we then proceed with the placement of item $x_{2}$. Its first agent, $p_{2}^{1}\in P$, will be placed a distance of $2$ to the right of $p_{1}^{s_{1}}$. Then $p_{2}^{2}$ is placed at a distance of $1$ to the right of $p_{2}^{1}$, $p_{2}^{3}$ is placed at a distance of $1+\varepsilon$ to the right of $p_{2}^{2}$, and so on. We repeat this process for all $x_{i}\in X$, in increasing subscript order. The seat graph $G$ consists of $K$ paths with $C$ vertices each, labelled $V^{1}$ to $V^{K}$. Figure 2 below shows an example of how two items from the Bin Packing are transformed into agents of the seat arrangement instance $I$.
[…]
Figure 2: Location of the agents transformed from $x_{1}$ and $x_{2}$ in the $1$-dimensional space, with $s_{1}=4$ and $s_{2}=2$
Now suppose that there is a packing of the items into $K$ bins. For each $\ell$ ($1\leq\ell\leq K$), suppose that the $\ell$-th bin contains items $x_{j_{1}},\dots,x_{j_{r}}$ for some $r\geq 1$, where $j_{1}<j_{2}<\dots<j_{r}$. Place the agents in $P_{j_{1}},P_{j_{2}},\dots,P_{j_{r}}$ onto the path $V^{\ell}$ in left-to-right order that respects the placement of these agents on the real line as constructed above. As every item $x_{i}$ is packed into a bin, all agents from $P_{i}$ are assigned a vertex in $G$ such that they feel no envy towards any other agent.
Conversely, suppose that there is an envy-free arrangement of the agents onto $G$. Let $k$ ($1\leq k\leq K$) be given and suppose that path $V^{k}$ contains an agent $p_{i}^{r}$. Then by the envy-free property, $V^{k}$ contains every agent in $P_{i}$. It follows that the packing obtained by placing item $x_{i}$ into bin $C_{k}$ whenever the family $P_{i}$ is placed on path $V^{k}$ is a legal packing of all items into $K$ bins. Therefore, as Bin Packing is NP-complete, so is seat arrangement with B-utility, $1$-dimensional strict preferences, and a seat graph in which every connected component is a path with the same number of vertices. ∎
Likewise, finding an envy-free arrangement is NP-hard when the seat arrangement instance $I$ has utility type S-utility. By Bodlaender et al. [bodlaender_hedonic_2020], this result holds even for symmetric preferences, by Massand and Simon [massand_graphical_2019] the result holds even for binary preferences, and by Ceylan [ceylan_optimal_2022] the result holds even for symmetric and binary preferences. We show that NP-hardness holds even for symmetric and strict preferences.
@@
Envy-Free Arrangement under S-utility is NP-hard even for symmetric and strict preferences and a seat graph with bounded degree where every graph component is either a cycle or a star.
@@
To prove this result, we will adapt the reduction from Theorem 3. The agent set $P$, and the seat graph $H$, and preference functions are identical. Regarding preferences, we start with the preference functions from the proof of Theorem 3. We then define new preferences $f_{a}^{\prime}(b)$ for each pair of agents $a,b\in P$ using the preferences $f_{a}(b)$ defined in the proof of Theorem 3 and a set of mutually different infinitesimaly small real numbers $\varepsilon_{i}$ for $i=1,\dots,|P|^{2}$ as follows: for each pair of agents $a,b$ we choose a different $\varepsilon_{i}$ and set $f_{a}^{\prime}(b)=f_{a}(b)+\varepsilon_{i}$.
Just like in the previous case, if $G$ can be partitioned into triangles we can build an envy-free arrangement, and if there is an envy-free arrangement every agent has utility at least $1$, as per Claim 1 from the proof of Theorem 3, and Claim 2 from the same proof still applies, with minor modifications to its proof. ∎
Similarly, in an instance with W-utility, finding an envy-free arrangement or reporting that none exists is NP-hard, even with symmetric and binary preferences, or symmetric and strict preferences, as we show with the following results.
@@
Envy-Free Arrangement under W-utility is NP-hard even for symmetric and binary preferences and a seat graph with maximum degree $2$ where every graph component is a $2$-cycle or a $3$-cycle.